find two consecutive even integers such that ten times the larger minus seven times smaller equals two times the sum of the integer. find the integer

Answer by ankor@dixie-net.com(22740)

Two consecutive even integers, x, (x+2)
:
Write an equation for exactly what it says
:
two consecutive even integers such that ten times the larger minus seven times smaller equals two times the sum of the integer.
10(x+2) - 7x = 2(x + x + 2)
10x + 20 - 7x = 2(2x + 2)
10x - 7x = 4x + 4 - 20
3x - 4x = -16
-x = -16
therefore
x = 16 and 18 are the two integers
:
:
See if that checks out
"two consecutive even integers such that ten times the larger minus seven times smaller equals two times the sum of the integer."
10(18) - 7(16) = 2(16+18)
180 - 112 = 2(34)
68 = 68